# Supervised Fine-tuning for Cross-Lingual Detoxification

### v2

@@ 1 Introduction
The global deployment of multilingual large language models (m-LLMs) has revolutionized cross-lingual natural language processing, enabling translation, comprehension, and generation capabilities across diverse languages (Weidinger et al., 2021; Üstün et al., 2024; Morris et al., 2020; Khondaker et al., 2024; Costa-jussà et al., 2024). However, this widespread adoption has simultaneously raised critical concerns about their tendency to generate harmful content across languages and linguistic contexts (Bommasani et al., 2021; Neplenbroek et al., 2024a; Jain et al., 2024; Bespalov et al., 2023; Sap et al., 2019; Ferrara, 2023; Neplenbroek et al., 2024b).
For instance, Figure 1 reveals a concerning phenomenon: LLMs can generate toxic completions for prompts like “Perdón, que elite de mierda” (English: “Sorry, what a shitty elite”) (Kenton et al., 2021; Li et al., 2024; Dementieva et al., 2024a; Wang et al., 2024b). However, our work shows that if we detox the model in one language, the detoxification can be observed in another language (SpanishGerman). There are several challenges hindering multilingual detoxification, which presents complexities beyond monolingual approaches due to varying toxic expressions across languages, different syntactic structures, and data scarcity in low-resource languages (Kirk et al., 2021; Beniwal et al., 2024; Xu et al., 2023; Dementieva et al., 2025b; Villate-Castillo et al., 2024).
@@
Key Findings: Our findings show that: (1) linguistic properties such as morphological complexity and syntactic structures may influence this cross-lingual toxicity transfer in languages with similar scripts and proportions, (2) Models like aya-expanse-8b (Dang et al., 2024) and bloom-7b (Scao et al., 2022), trained on English instances (High-resource language), show poor generalization to structurally different languages such as Chinese and Hindi (Figure 2). Furthermore, (3) the detoxification effects also vary across samples from different toxicity distributions like offensive, illegal, and hate-speech (Dubey et al., 2024; Koh et al., 2024)).
@@ 3 Experiments
Problem Setting Let be a set of different languages. Each language is associated with a dataset containing pairs of toxic and non-toxic sentences written in language . Detoxification is the task of using toxic sentences from language to update a language model such that it assigns a low probability to toxic sentences across all languages. More details in

### v3

# Breaking mBad! Supervised Fine-tuning for Cross-Lingual Detoxification

@@ 1 Introduction
The global deployment of multilingual large language models (m-LLMs) has revolutionized cross-lingual natural language processing, enabling translation, comprehension, and generation capabilities across diverse languages (Weidinger et al., 2021; Üstün et al., 2024; Morris et al., 2020; Khondaker et al., 2024; Costa-jussà et al., 2024). However, this widespread adoption has simultaneously raised critical concerns about their tendency to generate harmful content across languages and linguistic contexts (Bommasani et al., 2021; Neplenbroek et al., 2024a; Jain et al., 2024; Bespalov et al., 2023; Sap et al., 2019; Ferrara, 2023; Neplenbroek et al., 2024b). For instance, Figure 1 reveals a concerning phenomenon: LLMs can generate toxic completions for prompts like “Perdón, que elite de mierda” (English: “Sorry, what a shitty elite”) (Kenton et al., 2021; Li et al., 2024; Dementieva et al., 2024a; Wang et al., 2024b). However, our work shows that if we detox the model in one language, the detoxification can be observed in another language (SpanishGerman). There are several challenges hindering multilingual detoxification, which presents complexities beyond monolingual approaches due to varying toxic expressions across languages, different syntactic structures, and data scarcity in low-resource languages (Kirk et al., 2021; Beniwal et al., 2024; Xu et al., 2023; Dementieva et al., 2025b; Villate-Castillo et al., 2024).
@@
Key Findings: Our findings show that: (1) linguistic properties such as morphological complexity and syntactic structures may influence this cross-lingual toxicity transfer in languages with similar scripts and proportions,
(2) Models like aya-expanse-8b (Dang et al., 2024) and bloom-7b (Scao et al., 2022), trained on English instances (High-resource language), show poor generalization to structurally different languages such as Chinese and Hindi (Figure 2). Furthermore, (3) the detoxification effects also vary across samples from different toxicity distributions like offensive, illegal, and hate-speech (Dubey et al., 2024; Koh et al., 2024)).
@@ 3 Experiments
Problem Setting Let be a set of different languages. Each language is associated with a dataset containing pairs of toxic and non-toxic sentences written in language . Detoxification is the task of using toxic sentences from language to update a language model such that it assigns a low probability to toxic sentences across all languages. More details in Section §A.1.